OpenStudy (lsdxmdmaddicthc):

Can someone walk me through this... Everytime I do it, I get an answer that isn't in the multiple choices...

OpenStudy (lsdxmdmaddicthc):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 11 feet by 11 feet by 3 inches for a patio if the concrete costs $63.00 per cubic yard?

OpenStudy (lsdxmdmaddicthc):

Multiple choice answers : A. $211.75 B. $70.58 C. $317.63 D. $1,905.75

OpenStudy (amistre64):

convert your measures into yard units

OpenStudy (lsdxmdmaddicthc):

I did.

OpenStudy (lsdxmdmaddicthc):

I'll show all my work.

OpenStudy (lsdxmdmaddicthc):

Volume of Rectangular Prism = whl 11 feet = 132 inches 132*132*3 = 52,272 cubic inches

OpenStudy (lsdxmdmaddicthc):

52,272/12 = 4,356 cubic feet

OpenStudy (lsdxmdmaddicthc):

4,356/3 = 1,452 cubic yards

OpenStudy (lsdxmdmaddicthc):

1,452 yards * 63$ = $91,476

OpenStudy (lsdxmdmaddicthc):

That's a freaking expensive patio.....

OpenStudy (amistre64):

11 feet = 11 (1/3) yards, since a foot is 1/3 of a yard 3 inches = 3 (1/12) feet = 3(1/12)(1/3) yards

OpenStudy (lsdxmdmaddicthc):

I just converted it all to inches, and then back to feet, and then to yards... Wouldn't that work??

OpenStudy (amistre64):

only if you do it right, id just work it all to yards to avoid so many chances for error to creep in

OpenStudy (lsdxmdmaddicthc):

I can't find any errors....Can you? I will work it all to yards...

OpenStudy (amistre64):

(11/3 * 11/3 * 1/12) cubic yards, times 63

OpenStudy (lsdxmdmaddicthc):

Why 1/12?

OpenStudy (amistre64):

63*169/108

OpenStudy (lsdxmdmaddicthc):

Even that answer is wrong....

OpenStudy (lsdxmdmaddicthc):

There MUST be a typo in this question.....I'm not this bad at math lol :P

OpenStudy (amistre64):

3(1/12)(1/3) = 1/12

OpenStudy (amistre64):

3 inches is 3/12 = .25 feet, and .25 of 1/3 of a yard is well, 1/12 of a yard

OpenStudy (lsdxmdmaddicthc):

Right

OpenStudy (lsdxmdmaddicthc):

But even the answer we would get from that is "wrong"

OpenStudy (lsdxmdmaddicthc):

I'm going to bring this to my teacher's attention....

OpenStudy (amistre64):

11*11 = 121 .... not 169 :)

OpenStudy (amistre64):

63*121/108 is better

OpenStudy (lsdxmdmaddicthc):

WOOO! Alright. $70.58... Certainly a huge difference from spending Over 91 thousand dollars. Hahahaha

OpenStudy (lsdxmdmaddicthc):

Thanks man have an awesome day!!! ;)

OpenStudy (amistre64):

yeah your conversions must have been in error, the more you try to convert back and forth the more room there is for error in my approach, i made an error and that was only a single step ...

OpenStudy (lsdxmdmaddicthc):

True that.... I don't know where my error was but it's all okay :)

OpenStudy (amistre64):

if we convert to inches 1 cubic yard = 3*3*3 cubic feet = 36*36*36 cubic inches 11*12 = 132 inches 132*132*3 52272 --------- = ------- = 1.12 36*36*36 46656 63(1.12) = 70.56

OpenStudy (amistre64):

you didnt use cubic dimension 52272 cubic inches is not converted by 12 cubic inches: 12 is a linear measure, not a cubic measure. 12^3 is what you should have used 52272 cubic inches = 52272/12^3 cubic feet there arent 3 feet in a cubic yard, there is 3^3 cubic feet in a cubic yard that was your errors